Find the measure of an angle if its supplement measures 30 degrees more than twice its complement. Really struggling with this for some reason. can anyone explain how to try and tackle this wordy beast?!
Answer by Alan3354(21581)

Find the measure of an angle if its supplement measures 30 degrees more than twice its complement.
--------------
A = the angle
supplement = 180 - A
complement = 90 - A
---------------
(180 - A) = 2*(90 - A) + 30
180-A = 180 - 2A + 30
-A = -2A + 30
A = 30